OpenStudy (anonymous):

can someone help me with a k12 math question on slopes please, I will give a medal

OpenStudy (maddy1251):

I can. I go there.

OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: -2/3 A. y-7=2/3(x-1) B.y+7=2/3(x+1) C.y-7=-2/3(x+1) D.y+7=-2/3(x-1)

OpenStudy (anonymous):

I just need to know how to do this, I need to pass this XD

OpenStudy (maddy1251):

Do you know the equation for point-slope?

OpenStudy (maddy1251):

Not a problem. :-)

OpenStudy (anonymous):

not exactly...

OpenStudy (anonymous):

any way of help will be appreciated thanks :)

OpenStudy (anonymous):

are you there? @Maddy1251 ?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

are you guys there?

OpenStudy (mathstudent55):

The point slope equation of a line is: \(y - y_1 = m(x - x_1)\) where the slope is \(m\), and the point is \((x_1, y_1)\) You were given slope = m = -2/3, and point (1, 7). Use the equation above and (1) replace m with your given slope, \(-\dfrac{2}{3} \), (2) replace \(x_1\) with 1, and (3) replace \(y_1\) with -7. \( y - y_1 = m(x - x_1) \) \(y - (-7) = -\dfrac{2}{3}(x - 1)\) What does y - (-7) simplify to?

OpenStudy (anonymous):

im sorry XD I got the answer, and i closed this question but thanks for the help anyway XD